OpenStudy (anonymous):

Calculate the slope that passes through the points (7, -4) and (4,8) I got 4/-3 but that's not one of the answers...

OpenStudy (anonymous):

how did you get -4/3? show the steps.

OpenStudy (anonymous):

let me see. you did 8-(-4) = 4 for the numerator, correct?

OpenStudy (anonymous):

I did the y2 - y1 / x2 - x1 8-4/4-7 4/-3 /= fraction bar

OpenStudy (anonymous):

that's how I did it, but it's not the answer :/

OpenStudy (anonymous):

what is -(-2)?

OpenStudy (anonymous):

there isn't a -2 in the equation...

OpenStudy (anonymous):

well? what is the negative of a negative number?

OpenStudy (anonymous):

I know. answer my question though.

OpenStudy (anonymous):

a positive?

OpenStudy (anonymous):

yes. so then 8-(-4) = ?

OpenStudy (anonymous):

I re-did the problem and got 4

OpenStudy (anonymous):

that is wrong.

OpenStudy (anonymous):

positive 4

OpenStudy (anonymous):

you did 8-(-4)/4-7 = 12/-3 what is 12 divided by -3? +4 or -4?

OpenStudy (anonymous):

positive 4

OpenStudy (anonymous):

no. a positive number divided by a negative number gives a negative number. the answer is -4

OpenStudy (anonymous):

I checked answer 4 on my online school and it said it was correct, so if your right the online thing isn't ^^' sorry, I'm not saying your wrong it's just what my thing said.

OpenStudy (anonymous):

your online thing is wrong. there is only one answer possible for this and it is -4. talk to your teacher about it. maybe you will get extra credits for pointing out the mistake. apply the formula. the answer of the formula can be the only correct answer. click this link. http://www.google.com/search?client=ubuntu&channel=fs&q=12%2F%28-3%29&ie=utf-8&oe=utf-8

OpenStudy (anonymous):

Okay, I will when she gets back from vacation. But thank you for pointing out the mistake in my work.

OpenStudy (anonymous):

no problem